# Convergent transformations of visual representation in brains and models

### v1

@@ 4.5 Brain–Model alignment
This procedure, standard in recent literature 51, 64, ensures that alignment is captured independently of model depth. In cases—such as for language models—where correlations may be systematically negative, this approach preserves the interpretability of dissimilar representational geometries: negative alignment values indicate that stimuli which are close in one space (e.g., visual cortex) are systematically distant in another (e.g., language model embedding)89.
@@
where layers are indexed from 0 (first layer) to $|\mathcal{L}_{m}|-1$ (last layer), so $d^{*}_{p,r,m}\in[0,1]$. This procedure is standard for mapping the correspondence between brain regions and network hierarchy 19, 51. To obtain a subject-level measure for a set of models (e.g., all vision or language models), both the peak alignment and normalized depth measures were averaged across models. Group-level measures were then obtained by averaging subject-level values across participants.
@@ 4.6 Shared Component Extraction
To identify the dominant dimensions underlying shared representational geometry across participants, we applied kernel multi-view canonical correlation analysis (KMCCA) 33. KMCCA was performed on the set of image trials shared by all participants in NSD, extracting a low-dimensional subspace that maximized correlation across participants’ RDMs for each hub.
We used the KMCCA implementation from mvlearn90. For each hub, we extracted the first two components for exploratory purposes, and assessed the influence of the first (dominant) component in control analyses. While our focus was on the dominant axis, further work is needed to interpret all statistically significant components. Image-level semantic annotations and category labels were derived from MS-COCO 81. For automated labeling of the scene-to-object gradient, we used Pixtral-12B 91 (see Fig. 5a).
Further details of the KMCCA procedure, semantic projections, and partial RSA controls are provided in Supplementary Section A.5.
@@
Functional-MRI data and image stimuli were obtained from the Natural Scenes Dataset (accessed via https://registry.opendata.aws/nsd), BOLD5000 (10.18112/openneuro.ds001499.v1.3.0 and 10.1184/R1/c.5325683), and THINGS-fMRI (10.25452/figshare.plus.c.6161151.v1 and 10.17605/osf.io/jum2f). Image metadata were extracted from MS-COCO (https://cocodataset.org). Model checkpoints analyzed in this work are available on Hugging Face (https://huggingface.co/collections/pablomm/convergent-transformations-6808f7de248fa9674acac588). All resources were publicly accessible in July 2025.
@@ Acknowledgments
We thank J. García-Arch, M. Domínguez-Orfila, D. Pacheco-Estefan, C. Baldassano, and E. Cámara for helpful comments and discussions, the creators of the Natural Scenes Dataset, THINGS, and BOLD5000 for making their data publicly available, and the teams who developed and released the models on Hugging Face. This work was supported by the Spanish Ministerio de Ciencia, Innovación y Universidades, which is part of Agencia Estatal de Investigación (AEI), through the project PID2022 - 140426NB (Co-funded by European Regional Development Fund. ERDF, a way to build Europe). We thank CERCA Programme/Generalitat de Catalunya for institutional support.
@@
A Supplementary analyses
A.1 Hemispheric Asymmetry and the Validity of Symmetric Analyses
Figure A1: Hemispheric comparison of representational alignment. (A–C) Comparison of representational alignment (RSA, Pearson’s $r$; $N=8$, NSD) computed independently for the left (x-axis) and right (y-axis) hemispheres: (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment. Each point is a cortical parcel, colored by macro-anatomical group. The diagonal line indicates equal alignment. Squares denote parcels with a statistically significant hemispheric difference (paired $t$-test, $p<0.05$, FDR-corrected). (D) Boxplots of alignment within principal hubs. Significant hemisphere effects are marked. (E) Correlation between cross-hemispheric and bilateral inter-subject alignment. Cross-hemispheric RSA (y-axis) was computed between a parcel in one subject’s left hemisphere and the corresponding parcel in another subject’s right hemisphere, compared to bilateral RSA (x-axis). The near-perfect correlation ($r=0.99$) shows that combining hemispheres preserves geometry.
We quantified hemispheric asymmetry in representational geometry by computing alignment measures separately for the left and right hemispheres using the HCP-MMP1.0 atlas31. For each cortical parcel, we assessed (i) inter-subject alignment, (ii) brain–vision model alignment, and (iii) brain–language model alignment, using the same Natural Scenes Dataset (NSD)23 responses as in the main analyses (see Fig. A1).
Across all modalities, a largely symmetric pattern was observed (Fig. A1a–c): the rank-order of parcels was highly similar between hemispheres (Spearman’s $r$: inter-subject $=0.93$, vision $=0.92$, language $=0.82$), although some significant asymmetries were present. Specifically, early visual cortex (V1–V4) showed stronger alignment in the left hemisphere, while alignment was higher in the right hemisphere for ventral and LOTC hub parcels (hMT+, LO, TPOJ). Vision- and language-model alignment followed similar trends.
Grouping parcels by functional hub (Fig. A1d) highlighted these asymmetries: left hemisphere alignment was higher in early visual cortex, while right hemisphere alignment was higher in ventral and LOTC hubs.
To determine whether these asymmetries reflected differences in representational content or only magnitude, we compared cross-hemispheric RSA (e.g., participant $p$’s left V1 vs subject $q$’s right V1) to bilateral RSA (subject $p$’s left+right V1 vs subject $q$’s left+right V1). These measures were almost perfectly correlated across parcels ($r=0.99$, Fig. A1e), with no parcel deviating from the regression line. This demonstrates that combining hemispheres increases overall alignment strength without distorting the underlying representational geometry—no region showed a unique lateralized axis absent in the contralateral hemisphere.
Together, these results justify our approach of using symmetric (hemisphere-combined) parcels for the main analyses: (i) combining ROIs across hemispheres maximizes statistical power for multivariate comparisons without introducing distortions in representational geometry; (ii) hemisphere-separated ROIs remain informative for quantifying lateralization effects. This flexible approach supports both detailed lateralization analysis and robust, large-scale cross-dataset RSA experiments with joined hemispheres.
@@ A.1 Hemispheric Asymmetry and the Validity of Symmetric Analyses
Figure A2: Repetition-matching strategies and their impact on inter-subject RSA. To compute inter-subject RSA, we compared different strategies for matching the three trial repetitions for each shared image between any pair of participants. (A) Schematic of unshifted (repetition 1→1, 2→2, 3→3) versus shifted (cyclic 1→2, 2→3, 3→1) trial matching with example RDMs. (B, C)Group-average cortical RSA maps under unshifted (B) and shifted (C) matching; both reveal the early-visual, ventral, and LOTC hubs. (D) Parcel-wise comparison of shifted vs. unshifted RSA values shows a uniform attenuation under shifted matching, with two linear regimes reflecting stronger prefrontal attenuation.
A.2 Repetition-Matching Strategies for Inter-Subject RSA
@@
The Natural Scenes Dataset (NSD) presents 10,000 unique images across up to 40 scanning sessions per participant (30,000 trials total). A subset of 1,000 images (3,000 trials) was shared by all eight participants and always appeared in the same trial positions—interleaved among participant-unique stimuli—so that each shared image was viewed under identical practice, fatigue, and session-context conditions. This locked structure raises the possibility that inter-subject correlations might reflect non-perceptual factors (e.g., trial timing or memory demands) rather than purely stimulus-driven geometry.
@@
To isolate the stimulus-driven component, we adopted a “shifted-repetition” matching strategy when computing inter-subject representational similarity (IS-RSA). Rather than pairing each repetition index $k\leftrightarrow k$ across participants (“unshifted”), we cyclically permute so that repetition $k$ in participant $p$ matches repetition $(k\bmod 3)+1$ in participant $q$ (i.e. $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$; see Fig. A2A). Because all shared trials occupy the same session slots, this shift preserves image identity while breaking any exact repetition-locked confounds and also allows direct comparison with within-subject analyses (which must compare different repetitions; see Supplementary Fig. D5). Due to symmetry across participant-pair comparisons, shifts of 1 and 2 yield identical group-average maps.
@@
Recomputing the group-average cortical RSA under both unshifted (Fig. A2B) and shifted (Fig. A2C) schemes recovers the same three hubs—early visual cortex, ventral occipito-temporal parcels, and LOTC. As expected, shifting is more conservative: alignment magnitudes are uniformly lower (Supplementary Fig. A2D), yet the relative rank-order of all parcels remains unchanged (Spearman’s $\rho=0.96$). Frontal parcels (e.g. FEF, inferior frontal gyrus) suffer proportionally larger attenuation under shifting than occipito-temporal hubs, suggesting that these frontal regions encode task- or repetition-related variance (e.g. decision or attentional control) that the shifted criterion minimizes. By contrast, the robust persistence of occipital and temporal alignments underscores their stable, stimulus-driven representational geometry.
@@
To illustrate how these differences manifest in network structure, Fig. A3 shows representational connectivity graphs for both matching strategies. We first identified parcels whose inter-subject RSA exceeded an absolute threshold of $r=0.05$ in either the shifted (shift = 1 or 2) or unshifted maps. We then constructed a reduced connectivity matrix among this common set of nodes and extracted a two-iteration minimum-spanning-tree backbone (using edge weights defined as $1-\mathrm{RSA}$; see main Methods). The core two-stream topology (Early Visual $\rightarrow$ Ventral hub and Early Visual $\rightarrow$ LOTC hub) is preserved under both schemes. However, only the unshifted graph highlights FEF as a high-centrality node, further suggesting that frontal regions carry shared task- or repetition-locked signals that are attenuated by the shifted criterion.
@@
While these frontal effects are exploratory—given their lower RSA magnitudes and sensitivity to thresholding—they point to important future work on how task and attention signals coexist with stimulus-driven representations. At the same time, they underscore the necessity of shifted-repetition controls and of complementary interpretability techniques to isolate which stimulus-content features drive convergence, enabling mechanistic hypotheses about information transmitted across cortical routes.

### v2

@@ 4.5 Brain–Model alignment
This procedure, standard in recent literature 79, 49, ensures that alignment is captured independently of model depth. In cases-such as for language models-where correlations may be systematically negative, this approach preserves the interpretability of dissimilar representational geometries: negative alignment values indicate that stimuli which are close in one space (e.g., visual cortex) are systematically distant in another (e.g., language model embedding)80.
@@
where layers are indexed from 0 (first layer) to $|\mathcal{L}_{m}|-1$ (last layer), so $d^{*}_{p,r,m}\in[0,1]$. This procedure is standard for mapping the correspondence between brain regions and network hierarchy 10, 79. To obtain a subject-level measure for a set of models (e.g., all vision or language models), both the peak alignment and normalized depth measures were averaged across models. Group-level measures were then obtained by averaging subject-level values across participants.
@@ 4.6 Shared Component Extraction
To identify the dominant dimensions underlying shared representational geometry across participants, we applied kernel multi-view canonical correlation analysis (KMCCA) 40. KMCCA was performed on the set of image trials shared by all participants in NSD, extracting a low-dimensional subspace that maximized correlation across participants’ RDMs for each hub.
We used the KMCCA implementation from mvlearn68. For each hub, we extracted the first two components for exploratory purposes, and assessed the influence of the first (dominant) component in control analyses. While our focus was on the dominant axis, further work is needed to interpret all statistically significant components. Image-level semantic annotations and category labels were derived from MS-COCO 59. For automated labeling of the scene-to-object gradient, we used Pixtral-12B 1 (see Fig. 5a).
Further details of the KMCCA procedure, semantic projections, and partial RSA controls are provided in Supplementary Note 7.
4.7 Statistics and Reproducibility
Statistical testing and analysis were carried out in Python 3.10 using standard libraries (SciPy, Statsmodels). No statistical methods were used to pre-determine sample size; sample sizes were determined by availability in the public datasets: Natural Scenes Dataset ($N=8$), BOLD5000 ($N=4$), and THINGS-fMRI ($N=3$). These sizes are standard for deep-sampling fMRI studies where individual-subject reliability is high. All subjects and sessions from these datasets were included in the analysis. To assess the statistical significance of RSA maps (e.g., Fig. 2a-c), a two-tailed permutation test ($n=10,000$) was used to determine if RSA values significantly differed from a null distribution. For all other NSD group-level results (Figs. 1–5), significance was assessed using two-tailed paired t-tests ($N=8$) unless otherwise specified. In all cases, $p$-values were corrected for multiple comparisons using the False Discovery Rate (FDR) or Bonferroni method, with significance defined as $p<0.05$.
@@ 4.7 Statistics and Reproducibility
Functional-MRI data and image stimuli were obtained from the Natural Scenes Dataset (accessed via https://registry.opendata.aws/nsd), BOLD5000 (10.18112/openneuro.ds001499.v1.3.0 and 10.1184/R1/c.5325683), and THINGS-fMRI (10.25452/figshare.plus.c.6161151.v1 and 10.17605/osf.io/jum2f). Image metadata were extracted from MS-COCO (https://cocodataset.org). Model checkpoints analyzed in this work are available on Hugging Face (https://huggingface.co/collections/pablomm/convergent-transformations-6808f7de248fa9674acac588). Precomputed derivative data generated from these raw sources are available in Figshare 10.6084/m9.figshare.30753239. All resources were publicly accessible in January 2026.
@@ Acknowledgments
We thank J. García-Arch, M. Domínguez-Orfila, D. Pacheco-Estefan, C. Baldassano, and E. Cámara for helpful comments and discussions. This work was supported by the Spanish Ministerio de Ciencia, Innovación y Universidades, which is part of Agencia Estatal de Investigación (AEI), through the project PID2022 - 140426NB (Co-funded by European Regional Development Fund. ERDF, a way to build Europe). We thank CERCA Programme/Generalitat de Catalunya for institutional support.
@@
Competing interests
The authors declare no competing interests.
@@
Extended figures
Figure 1: Detailed parcel-level alignment for all modalities. Box plots show representational alignment scores computed for each of the 180 cortical parcels of the symmetric HCP atlas 34, organized by macro-anatomical groups 47. Each box represents the distribution of alignment scores across the eight participants ($N=8$). The red line and shaded area in each panel denote the mean and standard deviation of the null distribution, respectively, estimated via permutation testing. (A) Inter-subject alignment (IS-RSA). (B) Brain-to-vision-model alignment. (C) Brain-to-language-model alignment. This figure provides a detailed view of the results summarized in the main text, showing the parcel-by-parcel variability and confirming the concentration of high alignment within the Early Visual, Ventral, and LOTC hubs.
Figure 2: Cortical flat map projections of representational alignment. To provide a comprehensive view of their spatial distribution, group-level alignment scores are projected onto a flattened cortical surface. The figure shows (A) IS-RSA alignment, (B) vision-model alignment, and (C) language-model alignment. Values represent group-averaged RSA scores (Pearson’s $\rho$, $N=8$ from the NSD dataset), and major sulci are labeled for anatomical orientation. This visualization makes the full extent of the alignment patterns clear, particularly highlighting the widespread negative alignment (cool colors) between language models and the ventral visual stream, in contrast to the positive alignment seen for inter-subject and vision-model comparisons.
Figure 3: Comparison of within-subject and inter-subject representational analyses. (A) Cortical map of within-subject alignment (RSA, Pearson’s $r$). This was computed analogously to the inter-subject analysis by comparing RDMs from different trial repetitions within each participant, and then averaging across participants ($N=8$). The map reveals a similar spatial distribution to the inter-subject version, with a posterior-to-anterior gradient and three distinct hubs. (B) Parcel-wise comparison of within-subject and inter-subject alignment. The two measures are tightly correlated across the cortex, following a power-law relationship (fit shown in red; $R^{2}=0.98$). This indicates that brain regions that are consistent within an individual are also the ones that are consistent across individuals. (C) Within-subject representational connectivity matrix. Computed by correlating RDMs between all pairs of parcels within each participant, this analysis reveals the same three-hub block structure discovered using the inter-subject measurements. (D) Directed connectivity graph of the within-subject data, pruned using a 3-minimum spanning tree to highlight the network backbone. Edge color and width encode the inter-region alignment (RSA value), while node color represents each parcel’s peak alignment depth with vision models. Directionality is inferred from this depth measure based on a low-to-high sequential hierarchy, with arrows pointing from shallower- to deeper-aligning regions. The graph confirms the same information flow as the inter-subject analysis, with two streams emerging from early visual cortex: a medial-ventral stream to the Ventral hub and a lateral-dorsal stream to the LOTC hub.
Figure 4: Detailed layer-wise alignment profiles across cortical parcels. Detailed parcel-by-parcel view of the brain-model alignment curves summarized in the main text (Fig. 3), showing results for the 15 macro-anatomical groups with the highest inter-subject alignment. Only parcels that were themselves statistically significant in the inter-subject analysis are displayed. For each modality, the plotted alignment curves represent the RSA score averaged across all models of that type. Each individual line shows the mean alignment for a single parcel across participants ($N=8$), with shaded areas indicating the SEM. (A) Brain-to-vision-model alignment. This detailed view confirms that the distinct alignment profiles–decreasing, distributed, and increasing–are characteristic of the main hubs and often extend to anatomically adjacent parcels (e.g., the increasing profile of the LOTC hub is also observed in nearby STS regions). In contrast, areas with lower overall correspondence, such as in the prefrontal cortex, tend to show weak alignment across all model layers without a clear hierarchical preference. (B) Brain-to-language-model alignment. This view confirms that positive alignment is restricted to the LOTC hub and a small number of anatomically proximal parcels, all of which consistently exhibit the same step-like alignment profile.
Figure 5: Brain-model alignment profiles for different model families. Layer-wise alignment (RSA, Pearson’s $r$) is shown for representative parcels from the three main hubs, computed separately for each family of models (see Table 1 for a full list). Each curve represents the mean alignment across participants ($N=8$), with shaded areas indicating the SEM. Vision Models: Most vision model families (AugReg, CLIP, DINOv2) reproduce the characteristic hierarchical profiles: a decreasing alignment with Early Visual Cortex, a distributed alignment with the Ventral Hub, and an increasing alignment with the LOTC Hub. The Masked Autoencoder (MAE) models are a notable exception; consistent with their reconstructive training objective, they maintain a high alignment with Early Visual Cortex across all layers. Language Models: Despite differences in architecture, training data, and fine-tuning (e.g., instruction tuning in BLOOMZ), all language model families exhibit a similar pattern. Alignment is negligible or negative for Early Visual and Ventral hub parcels, while all families show the same characteristic step-function profile for parcels in the LOTC hub.
Figure 6: Noise-ceiling normalization of IS-RSA representational connectivity. (A) IS-RSA representational connectivity matrix (Pearson’s $\rho$) for the 30 parcels with highest inter-subject alignment. Black squares indicate, for each parcel (column), the parcel with which it shows maximal between-parcel IS-RSA connectivity. (B) Backbone graph derived from (A) using three iterations of a minimum-spanning-tree procedure. (C) Noise-ceiling–normalized connectivity matrix, obtained by dividing each edge by the geometric mean of the corresponding diagonal self-alignment estimates, setting the diagonal to 1. (D) Backbone graph extracted from the normalized matrix in (C). The preserved topology indicates that the inferred dual-route architecture is robust to parcel-wise differences in measurement reliability.
Figure 7: Whole-cortex representational connectivity network. Representational connectivity network among the 157 cortical parcels that exhibited statistically significant inter-subject alignment. The network is pruned using a 2-minimum spanning tree to highlight the strongest connections. To minimize confounds from anatomical variability, connectivity was computed on a within-subject basis ($N=8$); for each participant, we correlated the RDM of every parcel with every other parcel across different repetitions of the same stimuli, then averaged the resulting connectivity matrices. Edge color encodes the strength of this inter-region alignment (RSA), while edge width indicates the spanning tree order (1st or 2nd) to highlight the most central pathways. Node color represents each parcel’s peak alignment depth with vision models, providing a proxy for its hierarchical position. Arrows are added to the pruned edges to denote a statistically significant difference in this depth between connected parcels (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$), indicating the putative direction of information flow. The graph confirms the two primary visual processing streams identified in the main text. A medial-ventral stream connects early visual areas to core ventral stream regions involved in scene and object processing. In parallel, a lateral-dorsal stream connects early visual areas to the LOTC hub, including motion-sensitive areas (MT+ complex) and higher-order regions in the superior temporal sulcus (STS) and temporoparietal junction (TPOJ). The analysis also reveals key bridges between these streams, with nodes such as LO3 and PGp linking the ventral and LOTC hubs. Furthermore, despite weaker overall alignment, frontal and parietal regions like the Frontal Eye Fields (FEF) and the Medial Intraparietal Area (MIP) emerge as distinct hubs, suggesting their integration into this large-scale, stimulus-driven network.
Figure 8: Biological content selectively drives LOTC alignment and lateral‐stream connectivity. Scenes were split into those that contained biological agents (people or animals; 65.3% of the shared image set) and those that did not, and all inter-subject analyses were recomputed on each subset. (A) Parcel-wise inter-subject RSA within the three hubs. Blue: scenes with biological agents; orange: scenes without them. Early Visual parcels (V1–V4) show no reliable difference, Ventral parcels (VMV1–PHA3) align more for non-biological scenes, whereas LOTC parcels (V4t–TPOJ3) align almost exclusively for biological scenes. Stars mark two-tailed paired $t$-tests across eight participants ($^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$, FDR-corrected). (B,C) Group-level representational-connectivity matrices (left) and backbone graphs (right) for scenes without (B) and with (C) biological agents. Backbones were extracted with a three-iteration minimum-spanning-tree procedure. Node colour encodes each parcel’s peak vision-model layer (early $\rightarrow$ late: green $\rightarrow$ purple), node size scales with within-parcel IS-RSA, edge hue shows pairwise RSA, and edge width denotes the iteration (1–3) at which the edge entered the spanning tree. The lateral stream linking Early Visual Cortex to LOTC is largely absent for non-biological scenes (B) but re-emerges when biological agents are present (C), mirroring the parcel-level results in (A).
Figure 9: Image-based atlas of the dominant representational dimensions in Early Visual Cortex. The first two shared representational components for the Early Visual Cortex hub, derived using KMCCA on the responses of all participants to the shared image set. Each point is an individual stimulus image, and its colored border indicates its general semantic category (see legend). The layout reveals no clear clustering by semantic category. Instead, images that are close to each other in this shared space often have similar low-level visual properties (e.g., color palettes, textures, or global shapes), confirming that the geometry of this hub is primarily driven by visual similarity rather than abstract content.
Figure 10: Image-based atlas of the dominant representational dimensions in the Ventral Hub. The first two shared representational components for the Ventral hub, derived using KMCCA. Each point is an individual stimulus image. The layout reveals an organizing principle: a smooth gradient from scene-dominated images to object-dominated images. For example, the upper-left quadrant contains images of indoor spaces, the lower portion contains outdoor scenes, and the upper-right quadrant contains images where a single object is the primary focus. This evidences that the dominant representational axis of this hub recapitulates the classic scene-to-object gradient of ventromedial cortex.
Figure 11: Image-based atlas of the dominant representational dimensions in the LOTC Hub. The first two shared representational components for the LOTC hub, derived using KMCCA. Each point is an individual stimulus image, with its border colored by its general semantic category (see legend). The first KMCCA component (the x-axis) separates images containing biological agents (left cluster) from those without (right cluster). Within the biological cluster, there is a further separation, with images of animals generally located to the right of the images of people. This confirms that the LOTC hub’s geometry is primarily organized around the presence and type of animate entities in a scene.
Supplementary Notes
@@
Figure 12: Repetition-matching strategies and their impact on inter-subject RSA. To compute inter-subject RSA, we compared different strategies for matching the three trial repetitions for each shared image between any pair of participants. (A) Schematic of unshifted (repetition 1→1, 2→2, 3→3) versus shifted (cyclic 1→2, 2→3, 3→1) trial matching with example RDMs. (B, C)Group-average cortical RSA maps under unshifted (B) and shifted (C) matching; both reveal the early-visual, ventral, and LOTC hubs. (D) Parcel-wise comparison of shifted vs. unshifted RSA values shows a uniform attenuation under shifted matching, with two linear regimes reflecting stronger prefrontal attenuation.
Supplementary Note 1: Repetition-Matching Strategies for Inter-Subject RSA
@@
The Natural Scenes Dataset (NSD) presents 10,000 unique images across up to 40 scanning sessions per participant (30,000 trials total). A subset of 1,000 images (3,000 trials) was shared by all eight participants and always appeared in the same trial positions–interleaved among participant‐unique stimuli–so that each shared image was viewed under identical practice, fatigue, and session‐context conditions. This locked structure raises the possibility that inter‐subject correlations might reflect non‐perceptual factors (e.g., trial timing or memory demands) rather than purely stimulus‐driven geometry.
@@
To isolate the stimulus‐driven component, we adopted a “shifted‐repetition” matching strategy when computing inter‐subject representational similarity (IS‐RSA). Rather than pairing each repetition index $k\leftrightarrow k$ across participants (“unshifted”), we cyclically permute so that repetition $k$ in participant $p$ matches repetition $(k\bmod 3)+1$ in participant $q$ (i.e. $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$; see Fig. 12A). Because all shared trials occupy the same session slots, this shift preserves image identity while breaking any exact repetition‐locked confounds and also allows direct comparison with within‐subject analyses (which must compare different repetitions; see Supplementary Fig. 3). Due to symmetry across participant‐pair comparisons, shifts of 1 and 2 yield identical group‐average maps.
@@
Recomputing the group‐average cortical RSA under both unshifted (Fig. 12B) and shifted (Fig. 12C) schemes recovers the same three hubs–early visual cortex, ventral, and LOTC. As expected, shifting is more conservative: alignment magnitudes are uniformly lower (Supplementary Fig. 12D), yet the relative rank‐order of all parcels remains unchanged (Spearman’s $\rho=0.96$). Frontal parcels (e.g. FEF, inferior frontal gyrus) suffer proportionally larger attenuation under shifting than occipito‐temporal hubs, suggesting that these frontal regions encode task‐ or repetition‐related variance (e.g. decision or attentional control) that the shifted criterion minimizes. By contrast, the robust persistence of occipital and temporal alignments underscores their stable, stimulus‐driven representational geometry.
@@
To illustrate how these differences manifest in network structure, Fig. 13 shows representational connectivity graphs for both matching strategies. We first identified parcels whose inter-subject RSA exceeded an absolute threshold of $r=0.05$ in either the shifted (shift = 1 or 2) or unshifted maps. We then constructed a reduced connectivity matrix among this common set of nodes and extracted a two-iteration minimum‐spanning‐tree backbone (using edge weights defined as $1-\mathrm{RSA}$; see main Methods). The core two-stream topology (Early Visual $\rightarrow$ Ventral hub and Early Visual $\rightarrow$ LOTC hub) is preserved under both schemes. However, only the unshifted graph highlights FEF as a high-centrality node, further suggesting that frontal regions carry shared task- or repetition-locked signals that are attenuated by the shifted criterion.
@@
While these frontal effects are exploratory–given their lower RSA magnitudes and sensitivity to thresholding–they point to important future work on how task and attention signals coexist with stimulus‐driven representations.